OpenStudy (anonymous):

a spinner is divided into 10 equal sections numbered 1 through 10.if the arrow is spun,what is the probability the first number will be a 2 and the second number will be a 4? A. 0.05 B. 0.125 C. 0.2 D. 0.01

OpenStudy (anonymous):

BTW the arrow is spun twice

OpenStudy (anonymous):

to find this answer you will find the probability of landing on 2, then landing on 4. then you multiply the answers together.

OpenStudy (anonymous):

there are 10 spaces therefore the prob. of landing on two is 1/10. same goes for landing on 4.

OpenStudy (anonymous):

multiply both fractions together like so:\[\frac{ 1 }{ 10 }\times \frac{ 1 }{ 10 }\] which would give you \[\frac{ 1 }{ 100 }\] which is equal to.01

OpenStudy (anonymous):

This is an intersection probability. Since the outcomes are mutually exclusive, you multiply the two probabilities together to get the intersection probability.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

omfg u guys are so HELPFUL THANK U SOOO MUCH!!

OpenStudy (anonymous):

anytime friend!

OpenStudy (anonymous):

yes it does than k u so much

OpenStudy (anonymous):

you are very welcome

OpenStudy (anonymous):

i have another 1 for u.i f u have the time though?

OpenStudy (anonymous):

i have time for one more ;)

OpenStudy (anonymous):

you roll a die.find the probability of the number you end up rollong being a 3 or a 5. A. 1/3 B. 1/6 C.1/12 D. 1/2

OpenStudy (anonymous):

AND THAN U SOO MUCH U R VERY HELPFUL!

OpenStudy (anonymous):

no problem. in the future you can open a new question and then comment saying @montelbaun. it will message me with a link to the question. but for now, i will answer it here! one sec:

OpenStudy (anonymous):

o ok i didnt no u could do that. im still new to this just signed up yesterday:P

OpenStudy (anonymous):

okay: there are 6 sides to a die labeled 1-6. the probability of rolling a 3 is:\[\frac{ 1 }{ 6 }\] the prob. of rolling a 5 is also \[\frac{ 1 }{ 6 }\] you are only rolling the die once therefore, you will add the two fractions together.\[\frac{ 1 }{ 6 }+\frac{ 1 }{ 6 } =\frac{ 2 }{ 6 } or \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

and i am new too! just found this site about 3 days ago! its awesome!

OpenStudy (anonymous):

yea it is!. and u explain it so well thank u!

OpenStudy (anonymous):

no problem! did the process make sense?

OpenStudy (anonymous):

yes it did. i get it more now but wat if their was 2 dice?

OpenStudy (anonymous):

well it depends. if you want the probability of rolling a 3 or 5 on each die, you would multiply the answer i gave you by itself. as in: (1/3) x (1/3). giving you (1/9).

OpenStudy (anonymous):

if you just want the prob. of rolling a 3 on one and a 5 on the other. it would just be: (1/6)x(1/6). giving you (1/36)

OpenStudy (anonymous):

ohh ok so the diff. is that u multiply!? same as 3 dice? u would multiply the 3 out aswell? i get it now!

OpenStudy (anonymous):

that is correct! and so on for 4, 5, or 6 dice

OpenStudy (anonymous):

ok thank u so much! u r the most helpful so far!if i could give u a medal i would!

OpenStudy (anonymous):

looks like you already did for the last answer, that is good enough! glad i could help. ive got to go to work now but like i said, in the future put @montelbaun, and if I am on i will come to help!

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thank u very much!

OpenStudy (anonymous):

have a good one, cya!

OpenStudy (anonymous):

cya